sq.rt 3+4i+squ.rt 3-4i=????? please solve it as soon as possible.
 

Grade:12th pass

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Sqrt 3 + sqrt3 +4i-4i

=2( sqrt3)

Regards
Arun
4 years ago
 
Dear student
 
sqrt (3+ 4i ) = 2 + i
 
sqrt (3 – 4i) = i – 2
 
hence on adding = 2i
 
In case of any difficulty please feel free to ask.